OpenStudy (anonymous):

How can you tell before you divide 425 by 9 that the first digit of the quotient is in the tens place?

jimthompson5910 (jim_thompson5910):

Well assume that the first digit in the quotient is in the hundreds place so say the quotient is 100 (the smallest 3 digit number) which would mean 100*9 = 900

jimthompson5910 (jim_thompson5910):

that's too big, which means that the quotient has to be smaller than 100

OpenStudy (campbell_st):

well if you look at the 4... 9 doesn't divide into 4... so that means the answer can't be in the 100s so now look at 42 9 divides into 42 6 remainder 6... so the answer will be in the 60s... hope this helps

OpenStudy (anonymous):

Umm could you please dumb that down a bit?

jimthompson5910 (jim_thompson5910):

let's say we wanted to divide 360 by 9 360/9 = 40 40 is the quotient we could multiply both sides by 9 to get 360 = 9*40 notice how 360 equals 9 times the quotient of 40

jimthompson5910 (jim_thompson5910):

now let's say we go back to the original problem of 425/9 if x is the quotient, then 425/9 = x multiply both sides by 9 to get 425 = 9x

jimthompson5910 (jim_thompson5910):

Now if x is some 3 digit number, then let's assume that it's the smallest 3 digit number possible

jimthompson5910 (jim_thompson5910):

so let's assume that x = 100 if that were true, then 425 = 9x 425 = 9*100 425 = 900 but the right side is too big, so that means that x is too big which tells us that x is at most a 2 digit number

OpenStudy (anonymous):

Oh okay. thanks so much!

jimthompson5910 (jim_thompson5910):

yw